# Concept camera for the next-generation mm-wave cosmological surveys

### v1

@@ 3.1 Challenges
While there are no technological gaps that would prevent the realization of a spectroscopic survey camera as described, a successful implementation is expected to encounter a number of practical challenges nevertheless:
Stray-light and out-of-band radiation:
The camera concept is based on broad-band power detectors (such as KIDs), which are coupled to the incident radiation from a celestial source through the telescope and instrument optics, an antenna, and a network of lithographic transmission lines. However, the detectors may also respond to unintended sources of direct illumination. Because direct illumination may be broad band, in contrast to the 200 channels coupled via the antenna and filterbank, they must be suppressed 200 times below of what would be sufficient for a broad-band continuum camera. To avoid an undesired loss of sensitivity, the utmost care must be taken to avoid stray light and/or blue-leak contamination of the detectors, within the instrument, and in the telescope optical environment in which the camera will operate.
Standing waves:
@@
of the detectors may be necessary, both to keep power in the detectors localized long enough for effective translation into electronic signals, and to prevent cross-talk among the thousands of detectors residing on the same substrate. Given the number of detectors per pixel and overall, the cross-detector leakage of non-thermal phonons within a pixel needs to be suppressed at the 30 dB level, while across pixels it may need to be suppressed at the 50 dB to avoid an increase in the effective background noise levels due to cross-contamination.
Magnetic shielding:
@@
KIDs (and other superconducting detectors) can be sensitive to magnetic fields, especially during the cooldown phase, and therefore effective magnetic shielding of the detectors in the focal plane may be critical for achieving the nominal detector performance in a real-world operation.
Cryogenics:
@@
Achieving the necessary detector NEPs for background limited operation from the ground may require operation near or below 100 mK, although the non-thermal detection mechanism of KIDs may also allow operating them at a higher temperature, such as 260–280 mK without a significant performance degradation. Depending on the telescope it is operated on, the camera may have to be installed in the Cassegrain focus, in which case the dewar will tilt with telescope elevation, and pose additional challenges to maintaining cryogenic performance at the range of typical tilt angles during observations.
Multi-band capability:
If a multi-band instrument is desired with LIM as its primary science driver, all of the above will likely become significantly more challenging still. It is probably wisest to demonstrate and perfect a single-beam spectroscopic camera first, before taking the leap to an ambitious multi-band implementation.
Readout:
@@
The small pixel footprint will likely require readout multiplexing at the 10 GHz regime for KIDs. As such it may involve additional up- and downconverters to interface with digital-to-analog waveform outputs and the digitization of signals at a lower baseband.
@@
At the same time, a spectroscopic imager is exempt from some of the constraints that affect similar single-beam spectrometers, like SuperSpec. Notably, a multi-beam spectrometer does not require that all spectral channels are operational on any given pixel, or that channels are spaced at closely regular intervals in frequency, or that they provide closely uniform sensitivities. As the celestial source is scanned across many/all pixels during an observation[34], spectral information can be reconstructed effectively and uniformly as long as any given spectral ’channel’ is ’typically’ present, with a ’typical’ sensitivity, among the ensemble of pixels. As such, detector yields as low as 70–80% and/or gaps in the frequency coverage of individual pixels are generally tolerable, while non-uniformities will average out in the mapping process.
3.2 Scoping
@@
There is not one way to realize the spectroscopic imaging camera for the galaxy or LIM surveys. Depending on the availability of funding, resources, telescope, and the degree of ambition, there is ample room for adjusting the concept for specific needs or constraints. For example, one may scope (or descope) specific aspects of the camera concept:
Any number of bands (1–4), and any mm-wave band (850 m – 3 mm) can be scientifically interesting and useful – for deep CO galaxy surveys and LIM application alike.
Spectral resolution can be interesting and useful anywhere between 150 (2000 km/s) and 1000 (300 km/s). Low values may suffer from the dilution of the CO signal, when the line widths of the star-forming gas in the targeted galaxies do not fill the spectral channels. High values too may result in loss sensitivity if the galactic lines spread over multiple spectral bins. Optimal sensitivities are reached when the channel widths closely match the typical (mean) projected velocity dispersion of the star-forming gas. Assuming that 250 km/s projected velocity spreads are typical, 600 may be more optimal from a sensitivity perspective alone, but other constraints may push toward targeting different values.
@@
The camera’s field-of-view may be scaled to match what the host telescope optics can deliver. As the survey speed scales with the camera’s FoV, it is generally desirable to fill the available FoV with pixels to the maximal extent, if possible.
@@
The pixel packing can be adjusted. The spacing maximizes the survey speed for a fixed number of pixels[35], and relaxes the NEP requirement by focusing a full telescope beam on an individual pixel. However, smaller, more tightly packed pixels are also viable provided that the necessarily lower detector NEP values can be achieved for background-limited operation.
@@
It is not necessary for the camera to have dual-polarization sensitivity. A single-polarization instrument can do the same survey science. It just needs twice the time to reach equivalent sensitivities.
The number may be chosen to any convenient value for the implementation, and suitable re-imaging optics can interface the camera to the telescope, as appropriate.
@@ 4 IMPLEMENTATION
The proposed instrument is sufficiently complex that implementation within a single institution may prove challenging, as demonstrated by the most recent wave of large cameras with similar scope and complexity[36, 37, 25]. It is increasingly rare that a single institution would have all the resources (financial and human), as well as all the in-house expertise, necessary to cover every aspect of design, production, and fielding optimally; and to deliver flawless instrumentation on the typical timescales of funding cycles (3–5 years).
We therefore propose a new approach. It relies on the loose, open-source collaboration of partners, and benefits from the contribution of leading experts from all over the world. Participants who join shall generally agree to the following foundational principles:
@@
Participants agree to abide by a code of honor. (We are in this together because we all want to see the best spectrometer camera(s) be built and operated based on the concepts and designs we develop together.)
@@
Any participant in the collaboration may freely use, adapt, modify, or innovate any aspect of the camera concept in any way they deem appropriate, as long as they re-share all relevant new information (e.g. designs, simulations, lab measurements, documentation, publications) with the the rest of collaboration.
@@
These principles are meant to protect innovation while fostering a free-spirited collaboration among the partners. All information shared within the collaboration will likely be through a properly authenticated version control system (e.g. GitHub222https://www.github.com), such that once information is shared it remains accessible to all involved, but not to outside the collaboration. (The details of the version control system, organization, and workflows are to be identified later, and may evolve with the project.) Apart from sharing ideas, designs, and measurements, the collaboration will support its participants in any way possible to help them achieve their goals.
@@ 4.1 Modular Designs
While there is a uniting concept centered around a spectrometer camera capable of delivering the next-generation cosmological surveys of star-formation, it is not limited to the realization of a singular instrument or design that alone is capable of delivering on that goal. For every design choice one can make, alternative choices may be viable also. As such, our aim is to compile a small library of designs and components from which an instrument builder may draw to assemble a particular implementation of the camera concept. We may, therefore, have multiple alternatives for the detectors, focal-plane units, focusing elements, etc. And each module may in turn have sub-versions optimized somewhat differently, offering varying trade-offs. (Hopefully, common sense will guide us to converge on no more than a few alternative designs for the critical components).
@@ 5 CONCLUSION
We envision that a mm-wave spectroscopic imaging camera can be designed, developed, and built on a 5-year timescale, and deployed to a suitable 10-m class (or larger) ground-based mm-wave telescope to conduct the first truly volumetric surveys of star-formation in the Universe, in a dedicated effort over 1–3 years of operation. Such cameras can for the first time deliver a complete census of star-formation at 1–12, individually detecting tens of thousands of star-forming galaxies above a mass-luminosity threshold, complete with spectroscopic redshifts for each and every one. These cameras will be able to probe an order of magnitude deeper in star-forming galaxy masses and luminosities than any previous degree-scale mm-wave survey to date, and for the first time approach the galaxy populations that represent that bulk of the galaxies in the Universe. The same or similar camera can conduct Line Intensity Mapping (LIM) also, and the two complementary surveys can be conducted in a time-sharing arrangement for maximum scientific impact.
We believe that such a spectroscopic imaging camera, or family of cameras, are best built through a loose, open-source collaboration that involves a wide range of experts from around the globe, and in which participants are free to pursue their own specific goals and directions without being confined by the restrictive rules that are typical to consortia.
We hope that you, the reader of this paper, will join this collaboration to contribute your expertise to make these cameras the best they can be. If you are an instrument builder, we hope that you join because you can draw on the knowledge shared in the collaboration to build your capable instrument, or component, to its utmost potential and without delays and detours. And, we hope telescope organizations will join because they want a share of the transformative science that these cameras will deliver when deployed on a telescope they run.

### v2

@@ 4.1 Challenges
While there are no technological gaps that would prevent the realization of a IFU spectroscopic imaging survey instrument as described, a successful implementation is expected to encounter a number of practical challenges nevertheless:
Stray-light and out-of-band radiation. Our instrument concept is based on broad-band power detectors (such as KIDs), which are coupled to the incident radiation from a celestial source through the telescope and instrument optics, an antenna, and a network of lithographic transmission lines. However, the detectors may also respond to unintended sources of direct illumination. Because direct illumination may be broad band, in contrast to the 200 channels coupled via the antenna and filterbank, they must be suppressed 200 times below of what would be sufficient for a broad-band continuum camera. To avoid an undesired loss of sensitivity, the utmost care must be taken to avoid stray light and/or blue-leak contamination of the detectors, within the instrument, and in the telescope optical environment in which the instrument will operate.
@@
Phononic isolation of the detectors may be necessary, both to keep power in the detectors localized long enough for effective translation into electronic signals, and to prevent cross-talk among the thousands of detectors residing on the same substrate. Given the number of detectors per pixel and overall, the cross-detector leakage of non-thermal phonons within a pixel needs to be suppressed at the 30 dB level, while across pixels it may need to be suppressed at the 50 dB to avoid an increase in the effective background noise levels due to cross-contamination.
@@
Magnetic shielding. KIDs (and other superconducting detectors) can be sensitive to magnetic fields, especially during the cooldown phase, and therefore effective magnetic shielding of the detectors in the focal plane may be critical for achieving the nominal detector performance in a real-world operation.
@@
Cryogenics. Achieving the necessary detector NEPs for background limited operation from the ground may require operation near or below 100 mK, although the non-thermal detection mechanism of KIDs may also allow operating them at a higher temperature, such as 260–280 mK without a significant performance degradation. Depending on the telescope it is operated on, the IFU spectrometer may have to be installed in the Cassegrain focus, in which case the dewar will tilt with telescope elevation, and pose additional challenges to maintaining cryogenic performance at the range of typical tilt angles during observations.
@@
Readout. The small pixel footprint may require readout multiplexing at higher frequencies (e.g., 10 GHz) for KIDs, although the use of parallel-plate capacitors, instead of the more typical inter-digitated ones, can also shrink the size of resonators significantly[PRIMA-KIDS], provided these can be produced reliably and with sufficient targeting precision also. A higher frequency readout may involve additional up- and downconverters to interface with digital-to-analog waveform outputs and the digitization of signals at a lower baseband.
Operational challenges. While not inherent to the design or physical realization of the instrument per se, a common operational challenge for massively multiplexed KIDs is identifying which resonance belongs to which physical detector (i.e., spatial-spectral voxel). This may be addressed with suitable calibration hardware, such as wire-grid scanners in orthogonal directions[AMKID-new], and a sweeping LO for the frequency direction, together which can ’map’ resonators to voxels quickly. The readout position of resonators can be characterized and modeled in the lab.
None of the challenges listed above are show-stoppers. All of them can be (and have been) mitigated effectively, as have been demonstrated by the various on-chip spectrometer instruments and KID cameras already deployed or under development. However, failing to adequately address them all can lead to degraded instrument perfomance, and result in significantly (or vastly) reduced scientific yield. Thus, the optimal realization of such an instrument will likely require the expertise and contributions from multiple persons, groups, and/or institutions.
@@
At the same time, a spectroscopic imager is exempt from some of the constraints that affect similar single-beam spectrometers, like SuperSpec. Notably, a multi-beam spectrometer does not require that all spectral channels are operational on any given pixel, or that channels are spaced at closely regular intervals in frequency, or that they provide closely uniform sensitivities. As the celestial source is scanned across many/all pixels during an observation[Scanning], spectral information can be reconstructed effectively and uniformly as long as any given spectral ’channel’ is ’typically’ present, with a ’typical’ sensitivity, among the ensemble of pixels. As such, detector yields as low as 70–80% and/or gaps in the frequency coverage of individual pixels are generally tolerable, while non-uniformities will average out in the mapping process.
5 Scoping and Variants
@@
There is not one way to realize the IFU spectrometer for the galaxy or LIM surveys. Depending on the availability of funding, resources, telescope, preference for particular technologies, and the degree of ambition, there is ample room for adjusting the concept for specific needs or constraints. For example, one may scope (or descope) specific aspects of the instrument concept, or utilize alternative technologies to what we considered so far.
5.1 Telescope-specific variants
The instrument can be (and should be) optimized for the telescope it is ultimately deployed on, such as:
Larger telescopes (e.g. JCMT, IRAM 30-m[IRAM-30m], LMT[LMT], or AtLAST[AtLAST]) could do the same type of surveys faster and/or go deeper still. A 50-m class telescope may also provide astrometry of the sources at the 1–2′′ level, which is sufficient for direct follow-up in most other bands.
@@ 5.1 Telescope-specific variants
The IFU spectrometer field-of-view may be scaled to match what the host telescope optics can deliver. As the survey speed scales with the both the instrument’s FoV and with the telescope aperture size, it is generally desirable to fill the available FoV with pixels to the maximal extent possible, while larger telescopes, such as the IRAM 30-m, can deliver same results with a smaller FoV also.
5.2 Capabilities
There is flexibility on tuning the capabilities that the IFU spectrometer is designed to deliver, e.g.:
Any number of bands (1–4), and any mm-wave band (850 m – 3 mm) can be scientifically interesting and useful – for deep CO galaxy surveys and LIM application alike. Note, however, that for a multi-band instrument many of the technological challenges mentioned become aggravated.
Spectral resolution can be interesting and useful anywhere between 150 (2000 km/s) and 1000 (300 km/s). Low values may suffer from the dilution of the CO signal, when the line widths of the star-forming gas in the targeted galaxies do not fill the spectral channels. High values may be challenging to realize, and may also result in loss sensitivity if the galactic lines spread over multiple spectral bins. Optimal sensitivities are reached when the channel widths closely match the typical (mean) projected velocity dispersion of the star-forming gas. Assuming that 250 km/s projected velocity spreads are typical, 600 may be more optimal from a sensitivity perspective alone, but other constraints may push toward targeting different values.
5.3 Technologies
There are also technology driven choices that instrument builders can make. For example:
In this paper we focused on in-plane, on-chip spectrometers, but there are alternatives to achieving compactness, such as waveguide spectrometers[WSPEC] or vertical integration[Stover2024].
@@ 5.3 Technologies
The pixel packing can be adjusted. The spacing maximizes the survey speed for a fixed number of pixels[Holland2002], and relaxes the NEP requirement by focusing a full telescope beam on an individual pixel. However, smaller, more tightly packed pixels are also viable provided that the necessarily lower detector NEP values can be achieved for background-limited operation.
@@
Dual-polarization sensitivity may be achieved via suitable dual-polarization antenna designs[Bueno2018] also.
It is not necessary for the IFU spectrometer to have dual-polarization sensitivity. A single-polarization instrument can do the same survey science. It just needs twice the time to reach equivalent sensitivities.
The number may be chosen to any convenient value for the implementation, and suitable re-imaging optics can interface the instrument to the telescope, as appropriate.
@@ 6 Implementation
We initially conceived of building such an IFU spectrometer for the Greenland Telescope (GLT) specifically[GLT-survey], since the Center for Astrophysics Harvard & Smithsonian (CfA) is a partner institution of the GLT. Tim Norton was at the time the director of the GLT, while Thomas Greve is involved with the GLT from the Danish side. It has been clear to us from the outset that we need to collaborate with other groups and institutions to develop, build, and deploy such an instrument. It is also clear that other telescopes would be just as suitable platforms for the same type of instrument or science.
The proposed instrument is sufficiently complex that implementation within a single institution may prove challenging, as demonstrated by the most recent wave of large cameras with similar scope and complexity[SCUBA-2, TolTEC, AMKID]. It is increasingly rare that a single institution would have all the resources (financial, technological, and human), as well as all the in-house expertise, necessary to cover every aspect of design, production, and fielding optimally; and to deliver flawless instrumentation on the typical timescales of funding cycles (3–5 years).
We therefore propose a new approach. It relies on the loose, open-source collaboration of partners, and benefits from the contribution of leading experts from all over the world. For example, participants who join might generally agree to the following foundational principles:
@@
Participants agree to abide by a code of honor. (We are in this together because we all want to see the best IFU spectrometer(s) be built and operated based on the concepts and designs we develop together.)
@@
Any participant in the collaboration may freely use, adapt, modify, or innovate any aspect of the IFU spectrometer concept in any way they deem appropriate, as long as they re-share all relevant new information (e.g. designs, simulations, lab measurements, documentation, publications) with the the rest of collaboration.
@@
These principles are meant to protect innovation while fostering a free-spirited collaboration among the partners. All information shared within the collaboration will likely be through a properly authenticated version control system (e.g. GitHub111https://www.github.com), such that once information is shared it remains accessible to all involved, but not to outside the collaboration. (The details of the version control system, organization, and workflows are to be identified later, and may evolve with the project.) Apart from sharing ideas, designs, and measurements, the collaboration will support its participants in any way possible to help them achieve their goals.
The above listed ’rules’ of the collaboration are a starting point only. Ultimately, a collaboration, once formed, may decide a different set of rules, which promote progress and protect innovation in the way the participants feel is best for them.
@@ 6.1 Modular Designs
While there is a uniting concept centered around a IFU spectrometer capable of delivering the next-generation cosmological surveys of star-formation, it is not limited to the realization of a singular instrument or design that alone is capable of delivering on that goal. For every design choice one can make, alternative choices may be viable also. As such, our aim is to compile a small library of designs and components from which an instrument builder may draw to assemble a particular implementation of the IFU spectrometer concept. We may, therefore, have multiple alternatives for the detectors, focal-plane units, focusing elements, etc. And each module may in turn have sub-versions optimized somewhat differently, offering varying trade-offs. (Hopefully, common sense will guide us to converge on no more than a few alternative designs for the critical components).
@@ 7 Conclusion
We envision that a mm-wave IFU spectrometer can be designed, developed, and built on a 5-year timescale, and deployed to a suitable 10-m class (or larger) ground-based mm-wave telescope to conduct the first truly volumetric surveys of star-formation in the Universe, in a dedicated effort over 1–3 years of operation. Such IFU spectrometers can for the first time deliver a complete census of star-formation at 1–12, individually detecting tens of thousands of star-forming galaxies above a mass-luminosity threshold, complete with spectroscopic redshifts for each and every one. These types of instruments will be able to probe an order of magnitude deeper in star-forming galaxy masses and luminosities than any previous degree-scale mm-wave survey to date, and for the first time approach the galaxy populations that represent that bulk of the galaxies in the Universe. The same or similar instrument can conduct Line Intensity Mapping (LIM) also, and the two complementary surveys can be conducted in a time-sharing arrangement for maximum scientific impact.
We believe that such an IFU spectrometer, or family of instruments, are best built through a loose, open-source collaboration that involves a wide range of experts from around the globe, and in which participants are free to pursue their own specific goals and directions without being confined by the restrictive rules that are typical to consortia.
We hope that you, the reader of this paper, will join this collaboration to contribute your expertise to make these IFU spectrometers the best they can be. If you are an instrument builder, we hope that you join because you can draw on the knowledge shared in the collaboration to build your capable instrument, or component, to its utmost potential and without delays and detours. And, we hope telescope organizations will join because they want a share of the transformative science that these IFU spectrometers will deliver when deployed on a telescope they run.
@@
Disclosures
The authors declare that there are no financial interests, commercial affiliations, or other potential conflicts of interest that could have influenced the objectivity of this research or the writing of this paper.
Code, Data, and Materials Availability
All data in support of the findings of this paper are available within the article or as supplementary material.
Acknowledgements
AK, GK, and TN acknowledge support from the the Smithsonian Astrophysical Observatory.
TRG acknowledges financial support from the Cosmic Dawn Center (DAWN), funded by the Danish National Research Foundation (DNRF) under grant No. 140. TRG is grateful for support from the Carlsberg Foundation via grant No. CF20- 0534.
A previous version of this paper was published in the Proceedings of the SPIE Conference 13101[spie-paper].
Attila Kovács is a computer engineer at the Center for Astrophysics Harvard & Smithsonian. He received his A.B. degree in physics, astronomy and astrophysics from Harvard College in 1997, and his PhD degree in physics from the California Institute of Technology in 2006. His current research interests include submillimeter detector technologies and instrumentation, star-forming galaxies, signal processing, and data reduction and imaging algorithms.
Garrett K. Keating is an astrophysicist at the Center for Astrophysics Harvard & Smithsonian and is the Deputy Director of the Submillimeter Array. He received his B.A. in astrophysics from the University of California, Berkeley in 2008, and his Ph.D. in astrophysics from the University of California, Berkeley in 2016. His current research interests includes radio instrumentation, molecular gas in the early Universe, supermassive black holes, and atmospheric effects on interferometric observations.
Thomas R. Greve is a professor and co-director of the Cosmic Dawn Center, a center of excellence at the Technical University of Denmark and Copenhagen University since 2018. From 2012, Associate Professor at University College London. Obtained his PhD in 2005 from the Institute for Astronomy Edinburgh, followed by postdoctoral research positions at the California Institiute of Technology and the Max-Planck Institute for Astronomy. He led some of the first systematic surveys of molecular gas and dust in extremely star-forming galaxies in the young Universe. Since 2020, the Danish co-PI on the MIRI instrument on JWST, and the MIRI Deep Imaging Survey (MIDIS) of the Hubble Ultra Deep Field.
Timothy Norton is Director of the Submillimeter Array, Hawaii.
Biographies and photographs of the other authors are not available.
List of Figures
List of Tables